ENERGETICS OF THE GAS PHASE MONOHYDRATES OF TRANS-FORMANILIDE : A MICROSCOPIC APPROACH TO THE HYDRATION SITES OF THE PEPTIDE BOND

I. DIMICOLI, M. MONS, \underline{F. PIUZZI}, Laboratoire Francis Perrin, URA CEA-CNRS 2453, Centre d’Etudes de Saclay, Bat 522, 91191 Gif-sur-Yvette Cedex, France; and E.G. ROBERTSON, J.P. SI-MONS, Physical and Theoretical Chemistry Laboratory, University of Oxford, South Parks Road, Oxford OX1 3QZ, United Kingdom.

The energy of binding a water molecule to two H-bonded sites of trans-N-phenyl formamide (formanilide), an amide molecule that mimics the planar -NH-CO- peptide bond, has been measured in the gas-phase and calculated by an ab initio study. For the first time, the D0 dissociation energies of two 1:1 complexes, simultaneously observed in a supersonic expansion, is measured experimentally. They are found to be very similar. At the CO site, the water molecule acts as a proton donor and the H-bond has a energy of 5.40 +/- 0.28 kcal/mol. At the NH site, the water acts as an acceptor and the binding energy is 5.65 +/- 0.30 kcal/mol. Comparison of IR shifts with binding energies shows that they do not always correlate when the donor groups are different.